OpenStudy (anonymous):

The table below represents a linear function f(x) and the equation represents a function g(x): x f(x) -1 -9 0 -1 1 7 g(x) g(x) = 3x - 2 Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). (6 points) Part B: Which function has a greater y-intercept? Justify your answer. (4 points)

OpenStudy (anonymous):

dude we have the same question xD

OpenStudy (anonymous):

anybody?? T3T

OpenStudy (anonymous):

First function f(x)...\[slope = \frac{ y2-y1 }{x2-x1 }\] Let f(x) is the same as y. So waht is the slope?

OpenStudy (anonymous):

@phi @Compassionate @iPwnBunnies

OpenStudy (ipwnbunnies):

Use the slope formula Nelson gave you. Use two corresponding y-values and x-values on the table for f(x) to find the slope of that equation. Compare it to the slope of g(x).

OpenStudy (ipwnbunnies):

Also, the y-intercept is the value of the function when x = 0

OpenStudy (anonymous):

-9 and -1 for the Ys and 0 and -1 for the Xs?

OpenStudy (ipwnbunnies):

Yeah, make sure you match them with the subscripts though. (0, -1) is (X1,Y1) and (-1, -9) is (X2,Y2)

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

hmmm so then whats after that? i have this so far -> -9 and -1 for the Ys and 0 and -1 for the Xs for slope=y1-x1 over y2-x2.

OpenStudy (ipwnbunnies):

No, look again. \[slope = \frac{Y2 - Y1}{X2 - X1}\]

OpenStudy (anonymous):

-9 and -1 for the Ys and 0 and -1 for the Xs for slope=y2-y1 over x2-x1.

OpenStudy (ipwnbunnies):

Yes. Make sure they're in the right order, again! XD

OpenStudy (anonymous):

is that all i put for A?

OpenStudy (ipwnbunnies):

Yeah, that is part of A. And you'll also need to determine the slope of g(x). And compare the two slopes.

OpenStudy (anonymous):

oh crap >.<

OpenStudy (anonymous):

alright so how do i do that?

OpenStudy (ipwnbunnies):

How to do what? xD you didn't even find the slope for f(x) yet. Plus, function g(x) is written in slope-intercept form: \[y = mx + b\] Where 'm' is the slope of the function.

OpenStudy (anonymous):

yeah i meant how do i find the slope for -9-1 over 0-1

OpenStudy (ipwnbunnies):

Argh. You wrote it wrong again. >.< Ok. Let (X1, Y1) be the point (0, -1), as seen in the table (X2, Y2) is point (-1, -9) \[slope = \frac{-9 - (-1)}{-1 - 0} = \frac{-8}{-1} = 8\]

OpenStudy (anonymous):

ok got it so then is it f(x)=blank then compare that with g(x) = 3x - 2

OpenStudy (ipwnbunnies):

The slope of f(x) is 8, it is a line. And the slope of g(x) is 3. The bigger the slope, the steeper the line.

OpenStudy (anonymous):

and that would be the comparison correct?

OpenStudy (anonymous):

so then how bout for B)

OpenStudy (ipwnbunnies):

Mhm.

OpenStudy (ipwnbunnies):

For Part B, find the y-intercept for each function. Set x = 0, find the y-intercept. Compare them.

OpenStudy (anonymous):

the f(x) has a greater y-intercept because it has a steeper slope going from right to left?

OpenStudy (ipwnbunnies):

Nooo. The slope and y-intercept don't depends on each other. Just find the y-intercept of each function. .-. The higher y-intercept just means it crosses the y-axis at a higher value.